# Global regularity for solutions of magnetohydrodynamic equations with large initial data

### v7

@@ 1. Introduction
Then for each $L_{N}\geq 1$ there exist two positive numbers $c=c(N,\nu,\eta,L_{N})$ and $s_{5}=s_{5}(N,L_{N})$ such that
@@
$\int_{Q_{T}}\left[\left(\frac{1}{k_{n}^{\frac{\beta-1}{2}}}-\frac{1}{\varphi^{\frac{\beta-1}{2}}}\right)^{+}\right]^{\frac{4}{N}+2}dxdt$
$\geq$ $\int_{\{\varphi\geq k_{n+1}\}}\left[\left(\frac{1}{k_{n}^{\frac{\beta-1}{2}}}-\frac{1}{\varphi^{\frac{\beta-1}{2}}}\right)^{+}\right]^{\frac{4}{N}+2}dxdt$
@@
Moreover, the dependence of $c_{3}$ and $s_{0}$ on $r$ is such that both $\ln c_{3}$ and $s_{2}$ are smooth functions of $r$ over $(M_{q},\infty)$.
@@
(2.40) $\lim_{q\rightarrow\infty}M_{q}=\frac{N+2}{2}.$
@@
That is, $f(r)$ is convex over $(L_{N},\infty)$. We will exploit this property extensively. We may also assume that
$f^{\prime}(r)>0\ \ \mbox{for each $r>M_{q}$.}$
@@
$\leq$ $\max_{s\in[0,1]}|s^{r-\frac{N+2}{N}}\ln s|\int_{\{w<1\}}w^{\frac{N+2}{2}}dxdt\leq c.$
@@
This combined with (2.34) implies our theorem. That is to say, were (2.47) true for some $r>M_{q}$, we would have nothing more to prove.
We also need the following limit:
@@
$=$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}e^{\frac{-jf^{\prime}(r)}{\left(r-2j\right)}+\frac{f(r)}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}}.$
To continue, we refine our selection of parameters further. In view of (2.40), we may pick
(2.49) $q>M_{q}.$
@@
(2.50) $r>q.$
Introduce two new parameters
$A\in(0,\infty)\ \ \mbox{and}\ \ \xi\in(q,r).$
Consider the initial value problem
(2.53) $h^{\prime}(s)+\frac{h(s)}{s-M_{q}}$ $=$ $\left\{\begin{array}[]{ll}1&\mbox{if $s\in(q,\xi]$},\\ -A&\mbox{if $s\in(\xi,\infty)$,}\end{array}\right$.
(2.54) $h(q)$ $=$ $0.$
Solve this to obtain
$h(s)=\left\{\begin{array}[]{ll}\frac{s-M_{q}}{2}-\frac{(q-M_{q})^{2}}{2(s-M_{q})}&\mbox{if $s\in[q,\xi]$},\\ -\frac{A(s-M_{q})}{2}+\frac{(A+1)(\xi-M_{q})^{2}-(q-M_{q})^{2}}{2(s-M_{q})}&\mbox{if $s\in(\xi,\infty]$.}\end{array}\right$.
We can easily verify that $h(s)$ is Lipschitz on $[q,r]$. Furthermore, it is increasing on $(q,\xi)$ and decreasing on $(\xi,r)$. Hence,
(2.55) $h(s)>0\ \ \mbox{on $(q,r]$ whenever $h(r)>0$.}$
To verify the latter, we have
(2.56) $h(r)$ $=$ $-\frac{A(r-M_{q})}{2}+\frac{(A+1)(\xi-M_{q})^{2}-(q-M_{q})^{2}}{2(r-M_{q})}$
$=$ $\frac{-A\left[(r-M_{q})^{2}-(\xi-M_{q})^{2}\right]+(\xi-M_{q})^{2}-(q-M_{q})^{2}}{2(r-M_{q})}$
$=$ $\frac{-A(r-\xi)(r+\xi-2M_{q})+(\xi-q)(\xi+q-2M_{q})}{2(r-M_{q})}.$
Hence, to achieve the first inequality in (2.55), it is enough for us to take
(2.57) $A<\frac{(\xi-q)(\xi+q-2M_{q})}{(r-\xi)(r+\xi-2M_{q})}.$

### v8

@@ 1. Introduction
Then for each $L_{N}\geq 1$ there exist two positive numbers $c=c\left(N,\nu,\eta,L_{N}\right)$ and $s_{5}=s_{5}(N,L_{N})$ such that
@@
$\int_{Q_{T}}\left[\left(\frac{1}{k_{n}^{\frac{\beta-1}{2}}}-\frac{1}{\varphi^{\frac{\beta-1}{2}}}\right)^{+}\right]^{\frac{4}{N}+2}dxdt$ $\geq$ $\int_{\{\varphi\geq k_{n+1}\}}\left[\left(\frac{1}{k_{n}^{\frac{\beta-1}{2}}}-\frac{1}{\varphi^{\frac{\beta-1}{2}}}\right)^{+}\right]^{\frac{4}{N}+2}dxdt$
@@
Moreover, the dependence of $c_{3}$ and $s_{0}$ on $r$ is such that both $\ln c_{3}$ and $s_{0}$ are smooth functions of $r$ over $(M_{q},\infty)$.
@@
$\lim_{q\rightarrow\infty}M_{q}=\frac{N+2}{2}.$
@@
That is, $f(r)$ is convex over $(L_{N},\infty)$. We will exploit this property extensively. We also need the following limit:
Remark 2.1.
We may assume that
(2.46) $f^{\prime}(r)>0\ \ \mbox{for each $r>M_{q}$.}$
@@
$\leq$ $\max_{s\in[0,1]}|s^{r-\frac{N+2}{N}}\ln s|\int_{\{w<1\}}w^{\frac{N+2}{N}}dxdt\leq c.$
@@
This combined with (2.34) implies our theorem. That is to say, were (2.47) true for some $r>M_{q}$, we would have nothing more to prove. In this case, the constant in (1.5) also depends on $\|u^{(0)}\|_{2,\mathbb{R}^{N}}$ and $\|b^{(0)}\|_{2,\mathbb{R}^{N}}$. By the same token, we may also assume
(2.48) $f(s)>0\ \ \mbox{on $(M_{q},\infty)$.}$
However, neither (2.46) nor (2.48) will be used in the subsequent proof.
@@
$=$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}e^{\frac{-jf^{\prime}(r)+f(r)}{r-2j}}\|w\|_{q,Q_{T}}^{\frac{q(M_{q}-j)}{(q-L_{N})(r-2j)}}.$
To continue, we further refine our choice of parameters. We easily see from (2.18) and (2.33) that
$q-2M_{q}=q-\frac{2(q-2L_{N})}{\alpha q}=q-\frac{(N+2)(q-2L_{N})}{(q-N-2)}.$
By (2.41), $q-2M_{q}$ is a strictly increasing function of $q$ on $(N+2,\infty)$. Clearly, its range is $(-\infty,\infty)$. Hence, there is a unique $q_{0}\in(N+2,\infty)$ such that
$q_{0}=2M_{q_{0}}.$
Pick
(2.50) $q>q_{0},$
As a result,
$q>2M_{q}.$
@@
(2.51) $j>\frac{q}{2}.$
Introduce a new parameter
$\varepsilon>0.$
Then select
$r>2j+\varepsilon.$
This choice of $r$ satisfies (2.22). The introduction of $\varepsilon$ is to ensure that $r$ stays away from $2j$ because the constant $c$ and the exponents in (2.49) blow up as $r\rightarrow 2j$. In summary, we have
(2.52) $2M_{q}<q<2j<2j+\varepsilon<r.$
We can form the interpolation inequality
(2.53) $\|w\|_{2j+\varepsilon,Q_{T}}\leq\|w\|_{r,Q_{T}}^{\frac{r(2j+\varepsilon-q)}{(2j+\varepsilon)(r-q)}}\|w\|_{q,Q_{T}}^{\frac{q(r-2j-\varepsilon)}{(2j+\varepsilon)(r-q)}}$
On account of (2.52),
$j>M_{q}.$
We may raise both sides of (2.53) to the power of $\frac{(j-M_{q})(2j+\varepsilon)(r-q)}{(q-L_{N})(r-2j)(r-2j-\varepsilon)}$, thereby obtaining
$\|w\|_{2j+\varepsilon,Q_{T}}^{\frac{(j-M_{q})(2j+\varepsilon)(r-q)}{(q-L_{N})(r-2j)(r-2j-\varepsilon)}}\leq\|w\|_{r,Q_{T}}^{\frac{r(2j+\varepsilon-q)(j-M_{q})}{(q-L_{N})(r-2j)(r-2j-\varepsilon)}}\|w\|_{q,Q_{T}}^{\frac{q(j-M_{q})}{(q-L_{N})(r-2j)}}.$
Incorporating this into (2.49), we arrive at
(2.54) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}$
$+c\|w\|_{L_{N},Q_{T}}^{s_{1}}e^{\frac{-jf^{\prime}(r)+f(r)}{r-2j}+\frac{(2j+\varepsilon-q)(j-M_{q})f(r)}{(q-L_{N})(r-2j)(r-2j-\varepsilon)}}\|w\|_{2j+\varepsilon,Q_{T}}^{-\frac{(j-M_{q})(2j+\varepsilon)(r-q)}{(q-L_{N})(r-2j)(r-2j-\varepsilon)}}.$
Fix
$\eta\in(0,1).$
Without any loss of generality, we may assume
(2.55) $-jf^{\prime}(s)+f(s)+\frac{(2j+\varepsilon-q)(j-M_{q})f(s)}{(q-L_{N})(s-2j-\varepsilon)}$
$\geq$ $\left[\frac{(j-M_{q})(s-q)}{(q-L_{N})(s-2j-\varepsilon)}+\frac{(s-2j)(1-\eta)}{2j+\varepsilon-L_{N}}\right]f(2j+\varepsilon)\ \ \mbox{for each $s\in(2j+\varepsilon,r]$.}$
Indeed, suppose this is not true. That is, there is a $s\in(2j+\varepsilon,r]$ such that
$-jf^{\prime}(s)+f(s)+\frac{(2j+\varepsilon-q)(j-M_{q})f(s)}{(q-L_{N})(s-2j-\varepsilon)}$
$<$ $\left[\frac{(j-M_{q})(s-q)}{(q-L_{N})(s-2j-\varepsilon)}+\frac{(s-2j)(1-\eta)}{2j+\varepsilon-L_{N}}\right]f(2j+\varepsilon).$
Obviously, (2.54) holds for $r=s$. As a result, we can apply the preceding inequality to it. Upon doing so, we arrive at
(2.56) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}$
$+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{2j+\varepsilon,Q_{T}}^{-\frac{(j-M_{q})(2j+\varepsilon)(s-q)}{(q-L_{N})(s-2j)(s-2j-\varepsilon)}+\left[\frac{(j-M_{q})(s-q)}{(q-L_{N})(s-2j-\varepsilon)}+\frac{(s-2j)(1-\eta)}{2j+\varepsilon-L_{N}}\right]\frac{2j+\varepsilon}{s-2j}}$
$=$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{2j+\varepsilon,Q_{T}}^{\frac{(2j+\varepsilon)(1-\eta)}{2j+\varepsilon-L_{N}}}.$
As we noted earlier, $c$ and $s_{1}$ here remain bounded for $s\in(2j+\varepsilon,r]$. In view of (2.52), (2.16) and (2.41), we can form the interpolation inequality
$\|w\|_{2j+\varepsilon,Q_{T}}\leq\|w\|_{\infty,Q_{T}}^{\frac{2j+\varepsilon-L_{N}}{2j+\varepsilon}}\|w\|_{L_{N},Q_{T}}^{\frac{L_{N}}{2j+\varepsilon}}.$
Collect this in (2.56) to get
$\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}+\frac{L_{N}(1-\eta)}{2j+\varepsilon-L_{N}}}\|w\|_{\infty,Q_{T}}^{1-\eta}.$
Then (1.5) follows from a suitable application of Young’s inequality (1.10). That is to say, if (2.55) fails to be true, then (1.5) holds.
We next show that (1.5) remains valid under (2.55). On account of (2.45), the convexity of $f(s)$, there holds
(2.57) $f^{\prime}(s)\geq\frac{f(s)-f(2j+\varepsilon)}{s-2j-\varepsilon}\ \ \mbox{for each $s>2j+\varepsilon$}.$
We may decompose
$\frac{(j-M_{q})(s-q)}{(q-L_{N})(s-2j-\varepsilon)}=\frac{(j-M_{q})}{(q-L_{N})}+\frac{(j-M_{q})(2j+\varepsilon-q)}{(q-L_{N})(s-2j-\varepsilon)}.$
Incorporate the preceding two results into (2.55) to derive
(2.58) $-s_{j}f^{\prime}(s)+f(s)$ $\geq$ $\left[\frac{(j-M_{q})}{(q-L_{N})}+\frac{(s-2j)(1-\eta)}{2j+\varepsilon-L_{N}}\right]f(2j+\varepsilon)\ \ \mbox{for each $s\in(2j+\varepsilon,r]$.}$
where
(2.59) $s_{j}=j-\frac{(2j+\varepsilon-q)(j-M_{q})}{(q-L_{N})}.$
We further require
(2.60) $s_{j}>0.$
For the inequality to hold, it is enough for us to take
(2.61) $\varepsilon<\frac{q-L_{N}}{j-M_{q}}\left[j-\frac{(2j-q)(j-M_{q})}{q-L_{N}}\right].$
This is possible only when
(2.62) $\frac{(2j-q)(j-M_{q})}{q-L_{N}}-j=\frac{2j^{2}-(2q+2M_{q}-L_{N})j+M_{q}q}{q-L_{N}}<0.$
The numerator is a quadratic function in $j$. The discriminant is given by
$(2q+2M_{q}-L_{N})^{2}-8M_{q}q$ $=$ $4q^{2}+4M_{q}^{2}-4L_{N}(q+M_{q})+L_{N}^{2}$
$=$ $(2q-L_{N})^{2}+4M_{q}(M_{q}-L_{N})>0.$
The last step here is due to (2.42). According to the quadratic formula, the solution set to (2.62) is given by
$j\in(j_{1},j_{2}),$
where
$j_{1}$ $=$ $\frac{2q+2M_{q}-L_{N}-\sqrt{(2q+2M_{q}-L_{N})^{2}-8M_{q}q}}{4},$
$j_{1}$ $=$ $\frac{2q+2M_{q}-L_{N}+\sqrt{(2q+2M_{q}-L_{N})^{2}-8M_{q}q}}{4}.$
Obviously, both $j_{1}$ and $j_{2}$ are positive. In particular,
(2.63) $j_{2}\in(q,q+M_{q}-L_{N}).$
Under (2.60), we can write (2.58) in the form
$f^{\prime}(s)-\frac{f(s)}{s_{j}}$ $\leq$ $-\left[\frac{(j-M_{q})}{(q-L_{N})}+\frac{(s-2j)(1-\eta)}{2j+\varepsilon-L_{N}}\right]\frac{f(2j+\varepsilon)}{s_{j}}\ \ \mbox{for each $s\in(2j+\varepsilon,r]$.}$
Multiply through the inequality by $e^{-\frac{s}{s_{j}}}$ to derive
$\left(e^{-\frac{s}{s_{j}}}f(s)\right)^{\prime}\leq-\left[\frac{(j-M_{q})}{(q-L_{N})}+\frac{(s-2j)(1-\eta)}{2j+\varepsilon-L_{N}}\right]\frac{f(2j+\varepsilon)e^{-\frac{s}{s_{j}}}}{s_{j}}\ \ \mbox{for each $s\in(2j+\varepsilon,r]$.}$
Subsequently,
(2.64) $e^{-\frac{r}{s_{j}}}f(r)$ $=$ $e^{-\frac{2j+\varepsilon}{s_{j}}}f(2j+\varepsilon)+\int_{2j+\varepsilon}^{r}\left(e^{-\frac{s}{s_{j}}}f(s)\right)^{\prime}ds$
$\leq$ $e^{-\frac{2j+\varepsilon}{s_{j}}}f(2j+\varepsilon)+\frac{(j-M_{q})\left(e^{-\frac{r}{s_{j}}}-e^{-\frac{2j+\varepsilon}{s_{j}}}\right)f(2j+\varepsilon)}{(q-L_{N})}$
$+\frac{(1-\eta)f(2j+\varepsilon)}{2j+\varepsilon-L_{N}}\left[(r-2j)e^{-\frac{r}{s_{j}}}-\varepsilon e^{-\frac{2j+\varepsilon}{s_{j}}}+s_{j}\left(e^{-\frac{r}{s_{j}}}-e^{-\frac{2j+\varepsilon}{s_{j}}}\right)\right]$
$=$ $\left[1-\frac{j-M_{q}}{q-L_{N}}-\frac{(1-\eta)(\varepsilon+s_{j})}{2j+\varepsilon-L_{N}}\right]e^{-\frac{2j+\varepsilon}{s_{j}}}f(2j+\varepsilon)$
$+\left[\frac{j-M_{q}}{q-L_{N}}+\frac{(1-\eta)(r-2j+s_{j})}{2j+\varepsilon-L_{N}}\right]e^{-\frac{r}{s_{j}}}f(2j+\varepsilon).$
Recall (2.57) and (2.59) to obtain
(2.65) $-jf^{\prime}(r)+f(r)+\frac{(2j+\varepsilon-q)(j-M_{q})f(r)}{(q-L_{N})(r-2j-\varepsilon)}$
$\leq$ $-\frac{j(f(r)-f(2j+\varepsilon))}{r-2j-\varepsilon}+f(r)+\frac{(2j+\varepsilon-q)(j-M_{q})f(r)}{(q-L_{N})(r-2j-\varepsilon)}$
$=$ $\frac{(r-2j-\varepsilon-s_{j})f(r)}{r-2j-\varepsilon}+\frac{jf(2j+\varepsilon)}{r-2j-\varepsilon}.$
We further require
(2.66) $r>2j+\varepsilon+s_{j}.$
Under this choice for $r$, we can combine (2.64) and (2.65) to deduce
$-jf^{\prime}(r)+f(r)+\frac{(2j+\varepsilon-q)(j-M_{q})f(r)}{(q-L_{N})(r-2j-\varepsilon)}$
$\leq$ $\frac{(r-2j-\varepsilon-s_{j})}{r-2j-\varepsilon}\left[1-\frac{j-M_{q}}{q-L_{N}}-\frac{(1-\eta)(\varepsilon+s_{j})}{2j+\varepsilon-L_{N}}\right]e^{\frac{r-2j-\varepsilon}{s_{j}}}f(2j+\varepsilon)$
$+\frac{(r-2j-\varepsilon-s_{j})}{r-2j-\varepsilon}\left[\frac{j-M_{q}}{q-L_{N}}+\frac{(1-\eta)(r-2j+s_{j})}{2j+\varepsilon-L_{N}}\right]f(2j+\varepsilon)+\frac{jf(2j+\varepsilon)}{r-2j-\varepsilon}.$
Utilizing this in (2.54), we arrive at
(2.67) $\|w\|_{\infty,Q_{T}}$ $\leq$ $16\|w(\cdot,0)\|_{\infty,\mathbb{R}^{N}}+c\|w\|_{L_{N},Q_{T}}^{s_{1}}\|w\|_{2j+\varepsilon,Q_{T}}^{\beta_{2}},$
where
(2.68) $\beta_{2}$ $=$ $-\frac{(j-M_{q})(2j+\varepsilon)(r-q)}{(q-L_{N})(r-2j)(r-2j-\varepsilon)}+\frac{j(2j+\varepsilon)}{(r-2j-\varepsilon)(r-2j)}$
$+\frac{(r-2j-\varepsilon-s_{j})e^{\frac{r-2j-\varepsilon}{s_{j}}}(2j+\varepsilon)}{(r-2j-\varepsilon)(r-2j)}\left[1-\frac{j-M_{q}}{q-L_{N}}-\frac{(1-\eta)(\varepsilon+s_{j})}{2j+\varepsilon-L_{N}}\right]$
$+\frac{(r-2j-\varepsilon-s_{j})(2j+\varepsilon)}{(r-2j-\varepsilon)(r-2j)}\left[\frac{j-M_{q}}{q-L_{N}}+\frac{(1-\eta)(r-2j+s_{j})}{2j+\varepsilon-L_{N}}\right].$
We must have
(2.69) $\beta_{2}<\frac{2j+\varepsilon}{2j+\varepsilon-L_{N}}.$
Plug (2.68) into this and simplify the resulting inequality to obtain
$-\frac{(j-M_{q})(r-L_{N})}{(q-L_{N})}+r-M_{q}$
$+(r-2j-\varepsilon-s_{j})e^{\frac{r-2j-\varepsilon}{s_{j}}}\left[1-\frac{j-M_{q}}{q-L_{N}}-\frac{(1-\eta)(\varepsilon+s_{j})}{2j+\varepsilon-L_{N}}\right]$
$+(r-2j-\varepsilon-s_{j})\left[\frac{j-M_{q}}{q-L_{N}}+\frac{(1-\eta)(r-2j+s_{j})}{2j+\varepsilon-L_{N}}\right]<\frac{(r-2j)(r-L_{N})}{2j+\varepsilon-L_{N}},$
from whence it follows
(2.70) $\left[(\varepsilon+s_{j})e^{\frac{r-2j-\varepsilon}{s_{j}}}-(r-2j+s_{j})\right]\frac{(r-2j-\varepsilon-s_{j})\eta}{2j+\varepsilon-L_{N}}$
$<$ $\frac{(r-2j)(r-L_{N})}{2j+\varepsilon-L_{N}}+\frac{(j-M_{q})(r-L_{N})}{(q-L_{N})}-r+M_{q}$
$+(r-2j-\varepsilon-s_{j})\left[e^{\frac{r-2j-\varepsilon}{s_{j}}}\left[\frac{j-M_{q}}{q-L_{N}}+\frac{(\varepsilon+s_{j})}{2j+\varepsilon-L_{N}}-1\right]-\frac{j-M_{q}}{q-L_{N}}-\frac{(r-2j+s_{j})}{2j+\varepsilon-L_{N}}\right]$
$\equiv$ $h(r).$
We easily ckeck that
$\frac{(\varepsilon+s_{j})e^{\frac{r-2j-\varepsilon}{s_{j}}}}{2j+\varepsilon-L_{N}}-\frac{(r-2j+s_{j})}{2j+\varepsilon-L_{N}}>0\ \ \mbox{for $r>2j+\varepsilon$.}$
Indeed, there hold
$\left.(\varepsilon+s_{j})e^{\frac{r-2j-\varepsilon}{s_{j}}}-(r-2j+s_{j})\right|_{r=2j+\varepsilon}$ $=$ $0,$
$\left[(\varepsilon+s_{j})e^{\frac{r-2j-\varepsilon}{s_{j}}}-(r-2j+s_{j})\right]^{\prime}$ $>$ $0\ \ \mbox{for $r>2j+\varepsilon$.}$
Hence, we must show that there exists a $r>2j+\varepsilon+s_{j}$ such that
(2.71) $h(r)>0.$